OpenStudy (anonymous):

Given the polynomial f(x)=7x^2+13x-50, find f*|(0)

OpenStudy (anonymous):

I meant f(0).

OpenStudy (anonymous):

for that last part xD

OpenStudy (anonymous):

so everywhere you see an 'x', you replace that with 0: \[f(0) = 7(0)^{2}+13(0)-50\] since anything times 0 is 0, the first 2 terms equal 0 and we have: \[f(0) = 0+0-50 = -50\] so f(0) = -50